# MESA-QUEST: Tracing the formation of direct collapse black hole seeds via quasi-stars

### v2

@@ 1 Introduction
Most local galaxies appear to harbor a central supermassive black hole (SMBH). Dynamically determined masses of these central black holes reveal correlations with several key properties of their host bulges: masses, velocity dispersion and luminosities (L. Ferrarese & D. Merritt, 2000; S. Tremaine et al., 2002; J. Kormendy & L. C. Ho, 2013) suggesting that the SMBHs and the hosts co-evolve. Recent JWST observations show that many massive galaxies formed in the universe by $z\gtrsim 10$} (or $\lesssim 500$ Myr after the Big Bang), some populated with central black holes and many active galactic nuclei (AGN) are in place by $z\gtrsim 7$ (M. Castellano et al., 2022; R. P. Naidu et al., 2022; Y. Harikane et al., 2023; I. Labbé et al., 2023). Even further, observations suggest that actively accreting SMBHs with masses $\gtrsim 10^{7-8}$ M exist in high redshift galaxies such as UHZ1 (P. Natarajan et al., 2024). This early formation, in tandem with the galaxies’ high SMBH to stellar mass ratio of $\sim 10\%$, has been argued to provide evidence for the “direct collapse” scenario of SMBH formation (P. Natarajan et al., 2024; D. J. Whalen et al., 2023; E. R. Coughlin & M. C. Begelman, 2024; P. Natarajan et al., 2017). This mechanism predicts that heavy intermediate-mass black hole (IMBH) seeds form from the monolithic gravitational collapse of massive low-metallicity gas clouds in the cores of proto-galaxies (M. G. Haehnelt & M. J. Rees, 1993; A. Loeb & F. A. Rasio, 1994a; D. J. Eisenstein & A. Loeb, 1995; P. Madau & M. J. Rees, 2001; V. Bromm & A. Loeb, 2003a; M. C. Begelman et al., 2006; G. Lodato & P. Natarajan, 2006; T. Alexander & P. Natarajan, 2014; P. Natarajan et al., 2017).
@@
Given that BHs are expected to grow via mergers and accretion over cosmic time (M. G. Haehnelt et al., 1998; F. Pacucci & A. Loeb, 2020), information about their initial seeding is expected to be erased. Therefore, directly accessing the highest redshift population offers the best prospects (A. Ricarte & P. Natarajan, 2018; F. Pacucci & A. Loeb, 2022) to constrain BH seeding models. The origin of SMBHs in the Universe remains an open question at the frontier of astrophysics research. High-redshift data from JWST are providing new insights with the detection of a handful of $z>9$ quasars that lie close to the seeding epoch (see for e.g. R. Maiolino et al. (2023); A. Bogdan et al. (2023); A. D. Goulding et al. (2023); O. E. Kovács et al. (2024).
@@
The inferred masses of some of these earliest JWST SMBHs detected at $z>9$ are in the range of $10^{6-7}$ M; while those estimated using broad-line spectroscopy from the Sloan Digital Survey (SDSS), for example, are of the order of $\sim 10^{9}$ M at $z\sim 6-7$ (X. Fan et al., 2023). These estimates of SMBH mass have been used to place constraints on a combination of seeding and growth trajectories in these epochs (X. Fan et al., 2004; D. J. Mortlock et al., 2011; P. Natarajan, 2011; M. Volonteri, 2012; X.-B. Wu et al., 2015; E. Bañados et al., 2018).
@@
A range of theoretical seeding prescriptions that operate as early as $z\sim 20-25$ classified broadly as “light” and “heavy” seeding models have been proposed as starting points to account for the formation of observed SMBHs (P. Natarajan, 2011; M. Volonteri, 2012; T. E. Woods et al., 2019). Light seeds are believed to be remnants of the first generation of stars, the so-called Population III stars, that result in the production of initial BH seeds with 1010010-100 M (P. Madau & M. J. Rees, 2001).
@@
Heavy seed models, on the other hand, propose the formation of seeds $10^{4}\,-\,10^{5}\,M_{\odot}$ in several possible ways. First, heavy seeds could result from the direct collapse of pre-galactic gas disks (A. Loeb & F. A. Rasio, 1994b; M. Volonteri & M. J. Rees, 2005; G. Lodato & P. Natarajan, 2006; M. C. Begelman et al., 2006; G. Lodato & P. Natarajan, 2007). There are other promising proposed pathways and sites to form DCBH seeds at early epochs (i) from halos with supersonic baryon streaming motions relative to dark matter (A. Stacy et al., 2011; A. T. P. Schauer et al., 2017); (ii) in highly turbulent halos that are now thought to create the first quasars (M. A. Latif et al., 2022) and (iii) from major mergers of massive galaxies (L. Mayer et al., 2023; L. Mayer & S. Bonoli, 2019). An additional pathway involves rapid, amplified early growth of originally light seeds that may end up in conducive cosmic environments, such as gas-rich, dense nuclear star clusters (T. Alexander & P. Natarajan, 2014). Rapid mergers of light remnants in the early nuclear star clusters could also lead to the formation of heavy seeds at high redshifts as proposed by B. Devecchi & M. Volonteri (2009) as well as the runaway collapse of nuclear star clusters as proposed by M. B. Davies et al. (2011). In addition to these more conventional theoretical seeding models, primordial black holes (PBHs, S. Hawking 1971) that form in the infant Universe have also been explored as potential candidates to account for the origin of initial seeds for SMBHs (see N. Cappelluti et al. 2022 and references therein), as well as dark stars that are postulated to be powered by dark matter self-annihilation in their cores (K. Freese et al., 2016).
@@
The final stages of the formation of the DCBHs have been explored in simulations up to the stage where supermassive stars (SMSs) are produced. Rapid baryon collapse at the center of atomically cooled halos has been demonstrated to produce SMSs (T. Hosokawa et al., 2013; T. E. Woods et al., 2017; L. Haemmerlé et al., 2018; N. P. Herrington et al., 2023) which are then predicted to collapse by post-Newtonian GR instability or by depletion of core fuel at the end of post main sequence burning to form DCBHs via the formation of quasi-stars (M. C. Begelman et al., 2008). Simulations have yet to track the subsequent formation and growth of the BH seed an initio and the stellar component in the host galaxy.
@@
In this paper, we explore in detail, for the first time, the detailed evolution of the Quasi-Star stage using the MESA-QUEST code that was custom built for this purpose. The outline of our paper is as follows: in Sec. 2, we provide an outline of BH seed formation models and their cosmological contexts, and focus on tracing the DCBH formation process via quasi-stars in Sec. 3. Summarizing previous work on quasi-stars in Sec. 4, we then present our new implementation in MESA — which we call MESA-QUEST — with details that permit tracking their evolution and the results therein in Sec. 5. We focus on the key issue, the implementation of winds and mass loss in Sec. 6, and close with a discussion of our conclusions and prospects for future work.
2 Channels for BH seed formation
@@
The process by which supermassive black holes (SMBHs) form remains an open question. In principle, a stellar-mass black hole can reach $10^{9}$ M within the first Gyr after the Big Bang under the assumption that it accretes at the Eddington rate consistently for the entire time (e.g., Z. Haiman & A. Loeb, 2001). However, typical black hole formation through the collapse of a massive star releases large amounts of energy in the form of ultra-violet radiation, leaving the stellar remnant in a low-density environment and stifling its accretion rate for an extended period of time (M. A. Alvarez et al., 2009). This suggests that an alternate channel for the formation of SMBH seeds is necessary.
@@
The masses of the $z>9$ JWST SMBHs require heavy initial seeds, as argued for in the case of UHZ1 and GHZ9 (P. Natarajan et al., 2024; O. E. Kovács et al., 2024) and/or special growth conditions that allow extremely efficient and rapid accretion for GNz11 (R. Maiolino et al., 2024). In particular, the simultaneous detection of UHZ1 by JWST and the Chandra X-ray telescope; its IR colors; and the ratio of IR flux to X-ray match predictions and provide compelling evidence for a heavy initial BH seed formed via the direct collapse of gas (P. Natarajan et al., 2017).
@@
The key issue for the rapid growth of initial light BH seeds is the time crunch, which requires optimal growth conditions, such as abundant gas reservoirs to allow rapid accretion while suppressing feedback that could disrupt gas inflow (e.g., Z. Haiman & A. Loeb, 2001; F. Pacucci et al., 2015). Although these optimal conditions are more likely to be available at very early epochs, they are difficult to sustain continuously to grow seed BHs starting with masses of 10-100 M to the final SMBH masses powering $z>6$ quasars (K. Park & M. Ricotti, 2011; F. Pacucci et al., 2017; F. Pacucci & A. Loeb, 2020). This motivated the exploration of alternate seeding prescriptions, particularly those that might produce more massive initial seeds ranging in mass from $10^{3-5}$ M or those in which rapid amplification of accreted mass can be achieved by fine-tuning cosmic conditions (see M. Volonteri & J. Bellovary, 2012; Z. Haiman et al., 2013; P. Natarajan, 2014; T. E. Woods et al., 2019, for reviews).
@@
SMBH seeds are believed to have formed from pristine gas in the very early universe ($z\gtrsim 15$). Since atomic hydrogen is the only available coolant, the resulting Jeans masses can be large. Initial numerical studies of the formation of the first stars indicated that the initial mass function was likely tilted high compared to star formation occurring locally (V. Bromm & A. Loeb, 2003b; T. Abel et al., 2002). In more sophisticated recent simulations, fragmentation occurs readily, resulting in stars with lower mass than previously found (P. C. Clark et al., 2011; T. H. Greif et al., 2012; M. A. Latif et al., 2013; S. Hirano et al., 2014; A. Stacy et al., 2016). Notably, the remnants of this first generation of stars are expected to result in early light black hole seeds with masses ranging from 10-100 M. Although forming these light seeds from the remnants of the first stars seems natural in the early Universe, growing them to  $\sim\,10^{9}\,{M_{\odot}}$ in less than a Gyr, to account for the SDSS quasars, requires exquisitely fine-tuning conditions for gas supply and feeding. Light seeds have been shown to accrete gas with lower duty cycles and at sub-Eddington rates due to the typical cosmic environments that might harbor them (K. Inayoshi et al., 2016; F. Pacucci et al., 2017). Heavy seeds, meanwhile, produced from the direct collapse of pre-galactic gas disks in pristine halos were therefore proposed as an alternative path to account for the SMBHs powering the brightest and highest redshift quasars (G. Lodato & P. Natarajan, 2006; M. C. Begelman et al., 2006). In this DCBH seed formation picture, pristine gas in early disks that can become dynamically unstable will rapidly accrete mass to the center as their angular momentum can be rapidly dissipated through the formation of non-axisymmetric structures like bars on the small scales. This scenario, expected to operate in preferentially low angular momentum halos and hence low spin disks, would then in turn grow massive central objects (S. P. Oh & Z. Haiman, 2002; V. Bromm & A. Loeb, 2003a; G. Lodato & P. Natarajan, 2006; M. C. Begelman et al., 2006). DCBH formation occurs when the fragmentation of these gas disks is prevented by the suppression of cooling. This is achieved by ensuring rapid dissociation of any molecular hydrogen that might form, via irradiation of Lyman-Werner photons from an external field (C. Shang et al., 2010; B. Agarwal et al., 2013; J. A. Regan et al., 2017; E. Visbal & Z. Haiman, 2018).G. Lodato & P. Natarajan (2006, 2007) first demonstrated that connecting the larger scale cosmological context with the fate of these pristine gas disks could be used to derive global properties like the initial mass function for heavy DCBH seeds (G. Lodato & P. Natarajan, 2007). Radiative and hydrodynamical cosmological simulations are available to probe the availability of direct collapse seeding sites within the context of the standard paradigm for structure formation (B. Agarwal et al., 2013, 2014). These studies agree that seeding sites are common enough in the early universe to explain the existence of observed ($z\sim 6-7$) quasars, but maybe too rare to account for the existence of typical SMBH in a local $L_{*}$ galaxy (J. H. Wise et al., 2019; M. Habouzit et al., 2021). In addition to direct collapse of gas, heavy seeds can also form when an initially light seed under the right set of circumstances undergoes extremely rapid growth via super-Eddington accretion (in $\sim 10^{6}$ year or less) or when aided by mergers with other seeds (M. Volonteri & M. J. Rees, 2005; B. Devecchi & M. Volonteri, 2009; T. Alexander & P. Natarajan, 2014). Intermediate mass black hole seeds that result from the dynamical core collapse of the first star clusters could also be a channel for early seed formation (K. Omukai et al., 2008; B. Devecchi & M. Volonteri, 2009; N. C. Stone et al., 2017; P. Natarajan, 2020).
3 Tracking the evolution of Direct Collapse
@@
There are a number of possibilities and pathways for direct collapse to manifest, with the primary requirement being that typical fragmentation of the pre-galactic gas cloud must be prevented (i.e., temperatures $\gtrsim 10^{4}$ K and metallicities $\lesssim 10^{-5}~Z_{\odot}$) (M. C. Begelman & M. J. Rees, 1978; S. P. Oh & Z. Haiman, 2002). The most prominent coolant in these early proto-galactic gas clouds is molecular hydrogen ($\mathrm{H_{2}}$), which eventually causes this fragmentation. However, ionization by Lyman-Werner radiation (e.g., from a neighboring star-forming galaxy or from stripped binary stars, A. Ferrara & A. Loeb, 2013; M. Dijkstra et al., 2014; B. Agarwal et al., 2014; Götberg, Y. et al., 2019) can cause dissociation of $\mathrm{H_{2}}$ molecules and thus prevent fragmentation. Other possible direct collapse scenarios include the removal of angular momentum from non-axisymmetric perturbations and gravitational torques that occur during collapse, as well as after mergers of rare, high redshift massive galaxies when gas inflows cause an unstable gas disk to form and funnel in enough gas to collapse (M. C. Begelman & I. Shlosman, 2009; J.-H. Choi et al., 2013, 2015; D. Fiacconi & E. M. Rossi, 2015; L. Mayer et al., 2023).
@@
Another scenario results when nuclear fusion begins at the core of the collapsing gas cloud, forming a convective core surrounded by a convectively stable envelope in what is known as a supermassive star (M. C. Begelman, 2010). Eventually, the core collapses due to GR instability and a stellar-mass black hole is formed in the core. Alternatively, fusion may never occur, as it can be quenched by the continuous rapid infall of material (M. C. Begelman et al., 2006; E. R. Coughlin & M. C. Begelman, 2024). However, the heating and neutrino losses still result in the same fate: a stellar-mass black hole nested at the center of a massive, slowly rotating, accretion radiation-supported envelope of gas. Regardless of which of these processes occurs, this end result has become known as a Quasi-Star (M. C. Begelman et al., 2006, 2008; M. C. Begelman, 2010; W. H. Ball, 2012; D. Fiacconi & E. M. Rossi, 2015).
@@
Motivated by the need to account for many physical mechanisms to decipher the viability of quasi-stars to form heavy seeds, we developed MESA-QUEST to model and track the evolution of quasi-stars using MESA (B. Paxton et al., 2011, 2013, 2015, 2018, 2019) as it is capable of modeling them in greater detail with modern equations of state and opacities, detailed treatment of convection and winds, strong numerical stability, and with easier implementation of alternate physics. In this work, we begin with a description of previous work on quasi-stars, followed by a brief introduction to MESA. We then discuss our black hole implementation within MESA and explore both of the aforementioned boundary conditions. Finally, we explore multiple wind schemes, including the Reimers and “Dutch” schemes, which are built directly into MESA  in addition to implementing the custom schemes from C. Dotan et al. (2011) and D. Fiacconi & E. M. Rossi (2015).
These changes and any future work can be found in a publicly accessible GitHub repository. [^1] http://www.github.com/andysantarelli/MESA-QUEST
4 Previous Work on quasi-stars

### v3

@@ I Introduction
Most local galaxies appear to harbor a central supermassive black hole (SMBH). Dynamically determined masses of these central black holes reveal correlations with several key properties of their host bulges: masses, velocity dispersion and luminosities (FerrareseMerritt2000; Tremaine+2002; 2013ARA&A..51..511K) suggesting that the SMBHs and the hosts co-evolve. Recent JWST observations show that many massive galaxies formed in the universe by $z\gtrsim 10$} (or $\lesssim 500$ Myr after the Big Bang), some populated with central black holes and many active galactic nuclei (AGN) are in place by $z\gtrsim 7$ (Castellano_2022; Naidu_2022; Harikane_2023; Labb__2023). Even further, observations suggest that actively accreting SMBHs with masses $\gtrsim 10^{7-8}$ M exist in high redshift galaxies such as UHZ1 (Natarajan2024). This early formation, in tandem with the galaxies’ high SMBH to stellar mass ratio of $\sim 10\%$, has been argued to provide evidence for the “direct collapse” scenario of SMBH formation (Natarajan2024; Whalen2023; coughlin2024; Natarajan_2017). This mechanism predicts that heavy intermediate-mass black hole (IMBH) seeds form from the monolithic gravitational collapse of massive low-metallicity gas clouds in the cores of proto-galaxies (Haehnelt1993; loeb1994; eisenstein1995; Madau_2001; bromm&loeb2003; begelman2006; lodato2006; alexander2014; Natarajan_2017).
@@
Given that BHs are expected to grow via mergers and accretion over cosmic time (HaehneltPNRees1998; Pacucci_2020_growth), information about their initial seeding is expected to be erased. Therefore, directly accessing the highest redshift population offers the best prospects (RicartePN2018; Pacucci_Loeb_2022) to constrain BH seeding models. The origin of SMBHs in the Universe remains an open question at the frontier of astrophysics research. High-redshift data from JWST are providing new insights with the detection of a handful of $z>9$ quasars that lie close to the seeding epoch (see for e.g. Maiolino+2024; bogdan2023+; Goulding+2023; kovacs+2024.
@@
The inferred masses of some of these earliest JWST SMBHs detected at $z>9$ are in the range of $10^{6-7}$ M; while those estimated using broad-line spectroscopy from the Sloan Digital Survey (SDSS), for example, are of the order of $\sim 10^{9}$ M at $z\sim 6-7$ (fan+2023). These estimates of SMBH mass have been used to place constraints on a combination of seeding and growth trajectories in these epochs (Fan+2004; mortlock11; Natarajan2011; Volonteri_2012; Wu+2015; banados+2018).
@@
A range of theoretical seeding prescriptions that operate as early as $z\sim 20-25$ classified broadly as “light” and “heavy” seeding models have been proposed as starting points to account for the formation of observed SMBHs (Natarajan2011; Volonteri_2012; Woods_2018). Light seeds are believed to be remnants of the first generation of stars, the so-called Population III stars, that result in the production of initial BH seeds with 1010010-100 M (Madau_2001).
@@
Heavy seed models, on the other hand, propose the formation of seeds $10^{4}\,-\,10^{5}\,M_{\odot}$ in several possible ways. First, heavy seeds could result from the direct collapse of pre-galactic gas disks (loeb1994; volonteri05; lodato2006; begelman06; LodatoPN2007). There are other promising proposed pathways and sites to form DCBH seeds at early epochs (i) from halos with supersonic baryon streaming motions relative to dark matter (Stacy+2011; Schauer+2017); (ii) in highly turbulent halos that are now thought to create the first quasars (Latif+2022) and (iii) from major mergers of massive galaxies (Mayer+2023; Mayer_Bonoli2019). An additional pathway involves rapid, amplified early growth of originally light seeds that may end up in conducive cosmic environments, such as gas-rich, dense nuclear star clusters (AlexanderPN2014). Rapid mergers of light remnants in the early nuclear star clusters could also lead to the formation of heavy seeds at high redshifts as proposed by Devecchi_2009 as well as the runaway collapse of nuclear star clusters as proposed by Davies+2011. In addition to these more conventional theoretical seeding models, primordial black holes (PBHs, PBHs_1971) that form in the infant Universe have also been explored as potential candidates to account for the origin of initial seeds for SMBHs (see Cappelluti+2022 and references therein), as well as dark stars that are postulated to be powered by dark matter self-annihilation in their cores (Freese+2016).
@@
The final stages of the formation of the DCBHs have been explored in simulations up to the stage where supermassive stars (SMSs) are produced. Rapid baryon collapse at the center of atomically cooled halos has been demonstrated to produce SMSs (Hosokawa+2013; Woods+2017; Haemmerle+2018; Herrington_2023) which are then predicted to collapse by post-Newtonian GR instability or by depletion of core fuel at the end of post main sequence burning to form DCBHs via the formation of quasi-stars (Begelman_2008). Simulations have yet to track the subsequent formation and growth of the BH seed an initio and the stellar component in the host galaxy.
@@
In this paper, we explore in detail, for the first time, the detailed evolution of the Quasi-Star stage using the MESA-QUEST code that was custom built for this purpose. The outline of our paper is as follows: in Sec. II, we provide an outline of BH seed formation models and their cosmological contexts, and focus on tracing the DCBH formation process via quasi-stars in Sec. III. Summarizing previous work on quasi-stars in Sec. IV, we then present our new implementation in MESA — which we call MESA-QUEST — with details that permit tracking their evolution and the results therein in Sec. V. We focus on the key issue, the implementation of winds and mass loss in Sec. VI, and close with a discussion of our conclusions and prospects for future work.
II Channels for BH seed formation
@@
The process by which supermassive black holes (SMBHs) form remains an open question. In principle, a stellar-mass black hole can reach $10^{9}$ M within the first Gyr after the Big Bang under the assumption that it accretes at the Eddington rate consistently for the entire time (e.g., Haiman_2001). However, typical black hole formation through the collapse of a massive star releases large amounts of energy in the form of ultra-violet radiation, leaving the stellar remnant in a low-density environment and stifling its accretion rate for an extended period of time (Alvarez+2009). This suggests that an alternate channel for the formation of SMBH seeds is necessary.
@@
The masses of the $z>9$ JWST SMBHs require heavy initial seeds, as argued for in the case of UHZ1 and GHZ9 (Natarajan2024; kovacs+2024) and/or special growth conditions that allow extremely efficient and rapid accretion for GNz11 (Maiolino+2024). In particular, the simultaneous detection of UHZ1 by JWST and the Chandra X-ray telescope; its IR colors; and the ratio of IR flux to X-ray match predictions and provide compelling evidence for a heavy initial BH seed formed via the direct collapse of gas (Natarajan_2017).
@@
The key issue for the rapid growth of initial light BH seeds is the time crunch, which requires optimal growth conditions, such as abundant gas reservoirs to allow rapid accretion while suppressing feedback that could disrupt gas inflow (e.g., Haiman_2001; Pacucci_2015_growth_efficiency). Although these optimal conditions are more likely to be available at very early epochs, they are difficult to sustain continuously to grow seed BHs starting with masses of 10-100 M to the final SMBH masses powering $z>6$ quasars (Park+2011; Pacucci+2017; Pacucci_2020_growth). This motivated the exploration of alternate seeding prescriptions, particularly those that might produce more massive initial seeds ranging in mass from $10^{3-5}$ M or those in which rapid amplification of accreted mass can be achieved by fine-tuning cosmic conditions (see Volonteri&Bellovary2012; Haiman2013; Natarajan2014; Woods_2018, for reviews).
@@
SMBH seeds are believed to have formed from pristine gas in the very early universe ($z\gtrsim 15$). Since atomic hydrogen is the only available coolant, the resulting Jeans masses can be large. Initial numerical studies of the formation of the first stars indicated that the initial mass function was likely tilted high compared to star formation occurring locally (BL_03; Abel+2002). In more sophisticated recent simulations, fragmentation occurs readily, resulting in stars with lower mass than previously found (Clark+2011; Greif+2012; Latif+2013; Hirano+2014; Stacy+2016). Notably, the remnants of this first generation of stars are expected to result in early light black hole seeds with masses ranging from 10-100 M. Although forming these light seeds from the remnants of the first stars seems natural in the early Universe, growing them to  $\sim\,10^{9}\,{M_{\odot}}$ in less than a Gyr, to account for the SDSS quasars, requires exquisitely fine-tuning conditions for gas supply and feeding. Light seeds have been shown to accrete gas with lower duty cycles and at sub-Eddington rates due to the typical cosmic environments that might harbor them (Inayoshi+2016; Pacucci+2017). Heavy seeds, meanwhile, produced from the direct collapse of pre-galactic gas disks in pristine halos were therefore proposed as an alternative path to account for the SMBHs powering the brightest and highest redshift quasars (lodato2006; begelman06). In this DCBH seed formation picture, pristine gas in early disks that can become dynamically unstable will rapidly accrete mass to the center as their angular momentum can be rapidly dissipated through the formation of non-axisymmetric structures like bars on the small scales. This scenario, expected to operate in preferentially low angular momentum halos and hence low spin disks, would then in turn grow massive central objects (Oh_2002; bromm&loeb2003; lodato2006; begelman06). DCBH formation occurs when the fragmentation of these gas disks is prevented by the suppression of cooling. This is achieved by ensuring rapid dissociation of any molecular hydrogen that might form, via irradiation of Lyman-Werner photons from an external field (Shang+2010; Agarwal+2013; Regan+2017; Visbal_2018). lodato2006; LodatoPN2007 first demonstrated that connecting the larger scale cosmological context with the fate of these pristine gas disks could be used to derive global properties like the initial mass function for heavy DCBH seeds (LodatoPN2007). Radiative and hydrodynamical cosmological simulations are available to probe the availability of direct collapse seeding sites within the context of the standard paradigm for structure formation (Agarwal+2013; Agarwal+2014). These studies agree that seeding sites are common enough in the early universe to explain the existence of observed ($z\sim 6-7$) quasars, but maybe too rare to account for the existence of typical SMBH in a local $L_{*}$ galaxy (Wise+2019; Habouzit+2021). In addition to direct collapse of gas, heavy seeds can also form when an initially light seed under the right set of circumstances undergoes extremely rapid growth via super-Eddington accretion (in $\sim 10^{6}$ year or less) or when aided by mergers with other seeds (volonteri05; Devecchi_2009; AlexanderPN2014). Intermediate mass black hole seeds that result from the dynamical core collapse of the first star clusters could also be a channel for early seed formation (Omukai+2008; Devecchi_2009; Stone+2017; Natarajan2021).
III Tracking the evolution of Direct Collapse
@@
There are a number of possibilities and pathways for direct collapse to manifest, with the primary requirement being that typical fragmentation of the pre-galactic gas cloud must be prevented (i.e., temperatures $\gtrsim 10^{4}$ K and metallicities $\lesssim 10^{-5}~Z_{\odot}$) (begelman&rees1978; Oh_2002). The most prominent coolant in these early proto-galactic gas clouds is molecular hydrogen ($\mathrm{H_{2}}$), which eventually causes this fragmentation. However, ionization by Lyman-Werner radiation (e.g., from a neighboring star-forming galaxy or from stripped binary stars, Ferrara2013; Dijkstra2014; agarwal2014; Gotberg2019) can cause dissociation of $\mathrm{H_{2}}$ molecules and thus prevent fragmentation. Other possible direct collapse scenarios include the removal of angular momentum from non-axisymmetric perturbations and gravitational torques that occur during collapse, as well as after mergers of rare, high redshift massive galaxies when gas inflows cause an unstable gas disk to form and funnel in enough gas to collapse (Begelman_2009; Choi_2013; Choi2015; Fiacconi2015; Mayer+2023).
@@
Another scenario results when nuclear fusion begins at the core of the collapsing gas cloud, forming a convective core surrounded by a convectively stable envelope in what is known as a supermassive star (begelman2010). Eventually, the core collapses due to GR instability and a stellar-mass black hole is formed in the core. Alternatively, fusion may never occur, as it can be quenched by the continuous rapid infall of material (begelman2006; coughlin2024). However, the heating and neutrino losses still result in the same fate: a stellar-mass black hole nested at the center of a massive, slowly rotating, accretion radiation-supported envelope of gas. Regardless of which of these processes occurs, this end result has become known as a Quasi-Star (begelman2006; Begelman_2008; begelman2010; ball2012quasistars; Fiacconi2015).
@@
Motivated by the need to account for many physical mechanisms to decipher the viability of quasi-stars to form heavy seeds, we developed MESA-QUEST to model and track the evolution of quasi-stars using MESA (Paxton et al., 2011, 2013, 2015, 2018, 2019) as it is capable of modeling them in greater detail with modern equations of state and opacities, detailed treatment of convection and winds, strong numerical stability, and with easier implementation of alternate physics. In this work, we begin with a description of previous work on quasi-stars, followed by a brief introduction to MESA. We then discuss our black hole implementation within MESA and explore both of the aforementioned boundary conditions. Finally, we explore multiple wind schemes, including the Reimers and “Dutch” schemes, which are built directly into MESA  in addition to implementing the custom schemes from Dotan_2011 and Fiacconi2015.
These changes and any future work are publicly available on Zenodo with a living version available on GitHub. [^1] Zenodo: https://doi.org/10.5281/zenodo.18156852 (catalog 10.5281/zenodo.18156852)
GitHub: https://github.com/andysantarelli/MESA-QUEST
IV Previous Work on quasi-stars